OpenStudy (anonymous):

4/ 7x^-2/3 explain please

OpenStudy (anonymous):

\[\frac{4x^{-2/3}}{7}=\frac{4}{7\sqrt[3]{x^2}}\]

OpenStudy (anonymous):

not quite I'm afraid.. It's 4 over 7x^-2/3

OpenStudy (anonymous):

ohhhh no you got it